Chez has the following preferences over toothpaste, T, and a composite good, G, that represents all other goods. H (G, T ) = 100 × ln(T ) + G Chez receives $500 per month in income. The price of a tube of toothpaste, PT , is one dollar per tube and the price of the composite good is $.01 per unit of the composite good. (a) Determine the marginal utility of additional tube of toothpaste, the marginal utility of an additional unit of the composite good, and the rate at which Chez is willing to exchange toothpaste for the composite good?
Answer #1

Given H(G,T)=100*ln(T)+G

Marginal utility of additional tube of toothpaste=dH(G,T)/dT=MUT=100/T

Marginal utility of additional unit of composite good=dH(G,T)/dG=MUG=1

Rate at which Chez is willing to exchange toothpaste for composite good=MUG/MUT=T/100